OpenStudy (he66666):

linear algebra: linear transformations L: P2 -> P3 defined by L((p(t)) = t³p'(0) + t²p(0) (p'(t) means the derivative of p(t) with respect to t) How is this a linear transformation? I tried doing L[p(t) + q(t)] but it got so complicated since I got L[(t³p'(0) + t²p(0)) + (t³q'(0) + q²p(0))]..

OpenStudy (anonymous):

what is P2 and P3?

OpenStudy (he66666):

Sorry, it means polynomials

OpenStudy (anonymous):

hmm... let me see. Since the original polinomial is P2: for example p(t)=at^2 +bt +c the p'(t)=2at+b so L(p(t)) = t^3b + t^2c now: let q(t) = dt^2+et+f L(p(t) +q(t)) = L((a+d)t^2+(b+e)t +(b+c)) can you continue?

OpenStudy (he66666):

Oh, I think I got it!

OpenStudy (anonymous):

litle correction in L(p(t) +q(t)) = L((a+d)t^2+(b+e)t +(b+c)) should be L(p(t) +q(t)) = L((a+d)t^2+(b+e)t +(c+f)) then: L((a+d)t^2+(b+e)t +(c+f)) = t^3(b+e)+t^2(c+f)=t^3b +t^3e + t^2c+t^2f= =(t^3b+t^2c) + (t^3e+t^2f) = L(p(t)) +L(q(t))

OpenStudy (he66666):

Oh alright, thanks so much myko!

OpenStudy (anonymous):

yw